# MoM: Linear Sequence Modeling with Mixture-of-Memories

### v2

@@ Appendix E Training Loss Comparison
Figure 5: Training Loss. Loss curves for training 340M models on 15B tokens with a fixed random seed of 42.
Appendix F Fairness
@@
F.1 Equal Activated Parameter Count
@@
F.2 Equal Memory Capacity
@@
FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Number of Memories
Memories Activated
2 1
20.53 23.15 29.56 13.72 43.96 19.36 25.05
3 2
23.71 29.52 28.69 15.55 46.09 20.32 27.31
4 2
30.79 26.05 29.63 13.84 44.79 20.41 27.59
8 2
19.71 24.27 30.13 15.68 46.15 17.39 25.56
Shared Memory
w/ shared memory 30.79 26.05 29.63 13.84 44.79 20.41 27.59
w/o shared memory 25.61 27.09 28.35 14.32 41.94 19.02 26.06
Table 11: Complete Ablation Study Results. The complete results on recall-intensive tasks.
Wiki.
$\downarrow$
Lamb.
$\downarrow$
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Number of Memories
Memories Activated
2 1
26.84 66.26 46.25 23.21 35.54 25.40 64.69 51.30 41.07
3 2
25.86 52.78 45.62 23.46 36.25 28.84 65.13 53.67 42.16
4 2
26.00 51.25 46.13 24.15 35.91 28.26 65.61 52.57 42.10
8 2
26.40 52.30 45.66 24.74 35.93 28.39 64.91 49.49 41.52
Shared Memory
w/ shared memory 26.00 51.25 46.13 24.15 35.91 28.26 65.61 52.57 42.10
w/o shared memory 27.74 85.22 45.33 24.15 34.91 23.52 64.09 50.28 40.38
Table 12: Complete Ablation Study Results. The complete results on common-sense reasoning tasks.

### v3

@@ Appendix E Training Loss Comparison
Figure 7: Training Loss. Loss curves for training 380M models on 15B tokens with a fixed random seed of 42.
Appendix F The Hybrid of MoM and Transformer
Model FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16
MoM Hybrid 58.13 44.05 35.71 20.18 48.10 20.60 37.80
Model
ARC-e
[…]
ARC-c
$\mathrm{acc_{n}}\uparrow$
Hella.
$\mathrm{acc_{n}}\uparrow$
Lamb.
$\mathrm{acc}\uparrow$
PIQA
$\mathrm{acc}\uparrow$
Wino.
$\mathrm{acc}\uparrow$
Avg.
Transformer++ 44.91 25.94 34.95 26.90 64.31 51.07 41.35
MoM 44.65 24.74 36.54 27.93 66.16 51.78 41.97
MoM Hybrid 46.55 24.49 36.45 28.86 65.51 52.41 42.38
Appendix G Fairness
@@
G.1 Equal Activated Parameter Count
@@
G.2 Equal Memory Capacity
@@
Appendix H Detailed Scaling Results
To further examine the scalability of MoM from the perspective of memory capacity, we evaluate the effect of enlarging the memory pool beyond the main settings. Specifically, we compare two activation ratios, where the number of active memories accounts for either 0.5 or 0.25 of the total memory states. In all cases, an additional shared memory is included. Starting from a single memory as the baseline, we expand the number of memories to 2, 4, 8, and 16, and report the averaged results on both recall-intensive and commonsense benchmarks. The results, shown in Fig. 8, indicate that under a fixed activation ratio, increasing the memory size consistently improves performance on both categories of tasks.
Figure 8: Detailed scaling results. Performance shows a general improvement on both recall-intensive and commonsense tasks as the number of memories increases.
Appendix I Full Results
@@
Figure 9: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories without Aux Loss.